OpenStudy (anonymous):

16. Jesse played two days of golf. On the second day, he got a score of 2 below par, or -2. His total score for the two days was 8 below par or -8. What was Jesse's score the first day? a) Choose and label a variable for this situation. b) Write an equation using the variable that would answer the question. c) Solve your equation, and check your answer. d) Use the solution to write a sentence answer to the question in the word problem.

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

first day's score was 6 below par or -6

OpenStudy (anonymous):

two days score is 8 below par or, -8 second days score was 2 below par or, -2 First Day + Second Day = 8 below par or, -8 First Day + 2 below par or, -2 = 8 below par or, -8 First Day = -8 + 2 = -6

OpenStudy (lilg132):

total score for 2 days = -8 second day = -2 first day score = -8 + 2 first day score = -6

OpenStudy (lilg132):

b) Ts = total score f = first day score g = second day score ts = f + g g = ts - f

OpenStudy (lilg132):

or f = ts + g c) f = -8 + 2 f = -6

OpenStudy (anonymous):

sheg was these answered? b) Write an equation using the variable that would answer the question. c) Solve your equation, and check your answer. d) Use the solution to write a sentence answer to the question in the word problem.